OpenStudy (anonymous):

4. What is the slope of the line that passes through (3,2) and (5, 12)

OpenStudy (texaschic101):

There is a formula for this... slope = (y2 - y1) / (x2 - x1) now label your points....we can label our points as (x1,y1) and (x2,y2) (3,2)....x1 = 3 and y1 = 2 (5,12),,x2 = 5 and y2 = 12 now we just sub our info into the slope formula slope = (12 - 2) / (5 - 3) slope = 10/2 slope = 5 do you understand ? If you have any questions, please ask